OpenStudy (clara1223):

Given f(x)=-3x+(1/x)-3, simplify (f(x+h)-f(x))/h when x+4

OpenStudy (loser66):

Let me help you write it in neat \(f(x) = -3x+\dfrac{1}{x}-3 \) Simplify\( \dfrac{f(x+h)-f(x)}{h} \) when x = 4

OpenStudy (loser66):

Where are you stuck?

OpenStudy (clara1223):

@Loser66 I don't even know where to start...

OpenStudy (loser66):

from f(x), make the right hand side has the same denominator, what do you get?

OpenStudy (clara1223):

This is what I brainstormed so far, but I am not sure if this is anywhere close to the right track. Plus, it doesn't look like any of the answer choices

OpenStudy (clara1223):

@Loser66

OpenStudy (loser66):

Yes, but you make the same denominator wrong!! redo it here, please

OpenStudy (clara1223):

Which denominator are you talking about?

OpenStudy (loser66):

ok, let make it clear \(f(x) = -3x +\dfrac{1}{x}-3=\dfrac{-3x^2-3x+1}{x}\) ok?

OpenStudy (loser66):

Now, \(f(x+h) = \dfrac{-3(x+h)^2-3(x+h)+1}{x+h}\)

OpenStudy (loser66):

Open parentheses and Take difference \(f(x+h)-f(x) \) , we have \(\dfrac{-3(x^2+2xh+h^2)-3x-3h+1}{x+h}-\dfrac{-3x^2-3x+1}{x}\) right?

OpenStudy (loser66):

\(\dfrac{-3x^2-6xh-3h^2-3x-3h+1}{x+h}+\dfrac{3x^2+3x-1}{x}\) ok?

OpenStudy (loser66):

REMEMBER, that is just f(x+h)-f(x) , we didn't \(\div h\) yet!!!

OpenStudy (loser66):

Now, we have 2 fractions with different denominator; We need make them have the same denominator by multiply the first term by x, the second term by (x+h)

OpenStudy (loser66):

It becomes \(\dfrac{-3x^3-6x^2h-3xh^2-3x^2-3xh+x}{x(x+h)}+\dfrac{(x+h)(3x^2+3x+1)}{x(x+h)}\)

OpenStudy (loser66):

Simplify,

OpenStudy (loser66):

sorry, mistake at the last 1, it is -1, not 1 from numerator of the second term

OpenStudy (clara1223):

I understand, keep going please, you've been incredibly helpful.

OpenStudy (loser66):

just that, you simplify, then divided by h more, then replace x =4 to get the answer.

OpenStudy (loser66):

hey, barbecue, is there any shorter way?? Please, help. My way is tooooooooooooooo long and easy get mistake. hehehe.. barbecue @dan815 Please, please

OpenStudy (clara1223):

So far I have: \[\frac{ -3x ^{3}-6x ^{2}h-3xh ^{2}-3x ^{2}-3xh+x }{ x(x+h) }+\frac{ 3x ^{3}+3x ^{2}h+3x ^{2}+3xh-x-h }{ x(x+h) }\] after I canceled out: \[\frac{ 3x ^{2}h-3xh ^{2}-h }{ x(x+h) }\] Then I took h out of numerator: \[\frac{ h(3x ^{2}-3xh-1) }{ x ^{2}+xh }\] How am I doing so far?

OpenStudy (clara1223):

@Loser66 what's my next step?

OpenStudy (clara1223):

@dan815 can you help me? @loser66 was very helpful but he's right, his way is very long and complicated and I still am not clear on the subject.

OpenStudy (anonymous):

source: http://www.acalculator.com/math-calculators.html